40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 183 of 344  Not logged in ELOG logo
ID Date Author Type Category Subject
  8122   Wed Feb 20 20:58:37 2013 yutaUpdateAlignmentclipping centering checklist

Blue ones are the ones we checked yesterday.
Green ones are the ones we checked today.
Red ones are the ones we couldn't check.

We noticed mis-centering on green optics and partial clipping of higher order modes, but we did not touch any green optics in-vac. This is because green beam from Y end and X end has different spot positions on the green optics after periscopes. We confirmed that direct green beam from ends are not clipped.

I believe we have checked everything important. Any other concerns?

Attachment 1: ClippingCenteringChecklist.pdf
ClippingCenteringChecklist.pdf ClippingCenteringChecklist.pdf ClippingCenteringChecklist.pdf ClippingCenteringChecklist.pdf ClippingCenteringChecklist.pdf ClippingCenteringChecklist.pdf ClippingCenteringChecklist.pdf
  8121   Wed Feb 20 20:15:29 2013 yutaUpdateAlignmentSRM oplev status

Currently, SRM is misaligned in pitch so that SRM reflected beam hits on the top edge of SR3 (not on the mirror, but on the cage holding the mirror).
We also confirmed that SRM oplev beam is coming out from the chamber unclipped, and centered on QPD when SRM is "aligned".

  8120   Wed Feb 20 19:58:59 2013 ranaUpdateAlignmentBS table oplev re-arranged

Please confirm the SRM OL beam is not too bad and also find where the mis-aligned SRM puts its beam. WE want to be sure that there is not too much unwanted scattering from SRM into the PRFPMI.

  8119   Wed Feb 20 19:48:16 2013 yutaUpdateAlignmentBS table oplev re-arranged

[Sendhil, Yuta]

After aligning IFO and putting the access connector on, we also centered IPANG/IPPOS and all oplevs (except SRM).
To avoid clipping of PRM/BS oplevs, we re-arranged oplev steering mirrors on BS table.

What we did:
  1. Checked IPANG comes out unclipped after putting on the access connector.
  2. Centered IPANG on its QPD.
  3. Checked oplevs beams for ITMX/ITMY centered on in-vac mirrors, and centered them on their QPDs.
  4. Checked IPPOS beam is centered on the mirrors inside BS chamber, and centered IPPOS on its QPD.
  5. Tweaked oplev mirrors on BS chamber to make PRM/BS oplev beam unclipped and centered on mirrors, and centered them on their QPDs. To avoid clipping of oplev beams in BS table, we re-arranged oplev steering mirrors on BS table (outside the vaccum).


Current status:
  QPD values, IFO_ALIGN/MC_ALIGN screens, OSEM values attached.

  - IR incident beam and IFO aligned
  - X/Y end green coming out to PSL table (in higher order modes)
  - IPANG/IPPOS available
  - All oplevs available
  - AS/REFL/POP cameras ready
  - access connector, ETMX/ETMY heavy doors on
  - ITMX/ITMX/BS heavy doors are not on
  - AS/REFL/POP PDs not centered
  - POX/POY/TRX/TRY not aligned
  - AS beam coming out of the OMC chamber low by ~ 1 beam diameter (my bad)


Tomorrow:
  - Align AS/REFL/POP PD and lock PRMI
  - Take pictures of ITMX/ITMY/BS stacks
  - Put heavy doors on ITMX/ITMY/BS chambers
  - Start pumping down

Attachment 1: IFOALIGN_QPDs_OSEMs.png
IFOALIGN_QPDs_OSEMs.png
  8118   Wed Feb 20 19:20:50 2013 EvanUpdateAlignmentAS camera alignment

[Manasa, Evan]

Manasa and I are trying to get the AS beam onto the AS camera with a focusing lens. Currently, the mirror immediately preceding the camera has been removed and the camera and lens are sitting directly behind the BS.

  8117   Wed Feb 20 18:53:48 2013 JenneUpdateAlignmentIFO ready for doors, then pumping

[Yuta, Manasa, Sendhil, Jenne, Steve, Jamie, Koji, Evan]

The interferometer is well-aligned, and ready for pump-down.  The access connector is in place, as are the ETM heavy doors.  We will do ITM and BS doors tomorrow, then begin pumping.

Before we redid the ITM pointing, I confirmed that I could see both POX and POY on their respective tables, on a camera, unclipped.  I should check again quickly now that the ITM pointing has been finalized.

We went back to the arms, to perfect the ITM pointing.  Input beam was already centered at ETMY.  ETMY was pointing so that beam reflected to ITMY.  ITMY was adjusted a few (less than 4?) steps of 1e-3 size, to make reflected beam hit center of ETMY. 

BS was already pointing so beam hit center of ETMX.  ETMX was pointing to hit center of ITMX.  ITMX was adjusted a few (less than 4 again?) steps of 1e-3 size to make reflected beam hit center of ETMX.

Checked centering on AS path.  AS beam comes out of the vacuum a little low, but this wasn't discovered until after the access connector was in place.  We could adjust PZT3 (last AS mirror on BS table that sends beam over to OMC table), but we don't want to do this since we won't be able to re-confirm centering on the 3 mirrors on the OMC table.

Green beams (first Y, then X) were aligned using out-of-vac steering mirrors until beams were flashing in their respective arm cavities.  Green Y is a little close to the edge of the bottom periscope mirror, on the "up" periscope.  Since there is no steering between the arm and this periscope, fixing it would require moving the periscope.  We leave this to the next vent, when we finally install the BS table extension.  We were flashing a higher order yaw mode (5ish nodes) for the Y arm, and the very edge of the higher order mode on one side was a little bit clipped after reflecting off the steering mirror on the OMC table.  This is happening because that mirror is in the mount backwards (so we have access to the knobs).  We are confident that the straight-through beam is well centered on that mirror, so once we get it aligned to TEM00, there will be no clipping. We then did the X arm green, which was flashing a pitch higher order mode (again 5ish nodes).  The very edge of the higher order mode is clipping a little bit on the top mirror of the "down" periscope on the IMC table, but again the straight through beam is okay, and we're confident that the TEM00 mode will make it unclipped.  We could have touched some steering mirrors on the BS table, but since this was once upon a time well aligned, we don't want to futz with it.

Corner oplevs are all centered on their QPDs.  (The ETM oplevs were centered a few days ago).

Access connector and ETM doors are on.

The last 3 vertex doors will go on tomorrow when Steve gets in, and then we'll start pumping. 

There are no in-vac PZTs that need to be turned on (we've been using the output steering PZTs as non-energized fixed mirrors for some time), so we can lock at our leisure tomorrow afternoon.

  8116   Wed Feb 20 18:35:42 2013 JenneUpdateAlignmentNew in-vac alignment procedure

I have updated the vacuum checklist for in vacuum alignment.  Please take a look (https://wiki-40m.ligo.caltech.edu/vent/checklist) and see if I missed anything.  My goal was to make it incredibly step-by-step so there can be no mistakes.

  8115   Wed Feb 20 10:13:41 2013 yutaUpdateAlignmentPRC flashing brighter than last week

After in-vac alignment work last night, PRC is flashing brighter than PRMI alignment last week.
My hypothesis is that "we aligned PRM to junk MI fringe last week". Possibly, we used MI fringe caused by AR reflection of ITMs, or MI fringe reflected from SRM.

Videos:
  PRC flashing last week (youtube, elog #8085, elog #8091)

  PRC flashing this time (Lens in-front of AS camera was taken out)



My hypothesis can explain:
 - why we had dimmer POP last week (flash in half-PRC was way brighter even when we had more attenuators (youtube))
 - why I thought AS55 is broken (AS was too dim)

Conclusion:
  Be careful of junk beams.

  8114   Wed Feb 20 03:13:10 2013 yutaUpdateAlignmentclipping centering checklist

I attached clipping/centering checklist for the alignment.
Blue ones are the ones we checked today. Red ones should be checked tomorrow. Circles indicate centering on the optics, rectangles indicate clipping check, and arrows indicate retro-reflecting or bounces.
We found mis-centering on MMT1, PR2 and SR3 tonight (by ~0.5 beam diameter). They are also indicated.

I think we don't want to touch MMT1 and PR2 anymore, because they change input beam pointing.
I'm a little bit concerned about high beam on SR3, because we had PRC flashing in vertical higher order modes. We also see ETMX slider values high in pitch (~ 5.4).
Also, the diameter of ETMX reflected beam on ITMX looked larger and dimmer than ITMX transmitted beam, which doesn't seem reasonable.


Wednesday, Feb 20:
 - tweak TT1/TT2 and PRM so PRC flashes
 - re-check Yarm/Xarm bounces
 - center beam on all AS optics, starting from SR2
 - make sure REFL and AS is clear
 - check if TRY/TRX are coming out from the ends
 - check beam centering on mirrors in IMC/OMC chamber as far as you can reach
 - inject green from both ends
 - make sure green beams are not clipped by mirrors on BS chamber, IMC/OMC chamber
 - re-center all oplevs, with no clipping
 - check all OSEM values
 - take pictures of flipped PR2 and input TTs (and everything)
 - close all heavy doors and put the access connector back

Thursday, Feb 21:
 - make sure we can lock PRMI
 - start pumping down when Steve arrives

Attachment 1: ClippingCenteringChecklist.pdf
ClippingCenteringChecklist.pdf ClippingCenteringChecklist.pdf ClippingCenteringChecklist.pdf ClippingCenteringChecklist.pdf ClippingCenteringChecklist.pdf ClippingCenteringChecklist.pdf ClippingCenteringChecklist.pdf
  8113   Wed Feb 20 01:40:37 2013 ManasaUpdateAlignmentFinal IFO alignment- in progress

[Yuta, Sendhil, Jamie, Jenne, Rana]

1. After the MC centering, we tried to align the IFO using IPPOS and IPANG as reference. This did not recover the alignment perfectly. We were clipping at the BS aperture. Using TTs, we centered the beam at BS and PRM.
2. Using TTs, the beam was centered at ITMY and ETMY.
3. IPPOS and IPANG mirrors in-vacuum were aligned and were centered at the out-of-vacuum optics.
4. We checked the centering of the beam on optics in the BS and ITMY chamber. (Yuta will make an elog with the layout)
5. We retro-reflected ITMY at the BS and aligned ETMY such that we saw a couple of bounces in the arm cavity.
6. Using BS, the beam was steered to go through the center of ITMX and ETMX.
7. At this point we were able to see the MI fringes at the AS port.
8. We made fine alignments to the ITMX such that we saw MI reflected at the Farday.
9. We retro-reflected ITMX and aligned ETMX to see the beam bounce at the ITMX.
10. We aligned PRM such that PRC flashes. But we were not happy with the flashes (they were in higher order modes). We suspect that minor tuning of the input pointing might be necessary.
11. We closed for the day

  8112   Tue Feb 19 19:55:52 2013 yutaUpdateIOOMC yaw input tuned

[Jenne, Manasa, Yuta]

Since we levelled IMC stack, we had to center beam spots on MC mirrors again.
We did this by steering PSL mirror in yaw (about same amount but opposite direction to what we did in elog #8077)
Residual beam tilt compared with a line through MC1 and 3 actuator nodes is ~ 15 mrad, mainly in yaw.
MCdecenter_19Feb2013.png

  8111   Tue Feb 19 17:14:56 2013 ChloeUpdate QPD Circuit

I finished working out the circuit and figured out where the broken connections were. This is diagrammed in my notebook (will draw up more nicely and include in future elog post). Within the QPD circuitry, it seems like there are already opamps which regulate the circuit. I need to discuss the final diagram further with Eric.

I rewired the circuit inside the QPD box, which took awhile because it was difficult to solder the wires to such small locations without having multiple wires touch. This is completed, and on Friday I will begin to make the circuit to add/subtract signals to give pitch and yaw.

  8110   Tue Feb 19 15:40:34 2013 CharlesUpdateISSISS Prototype

After spending a good deal of time learning how to use the SR785, I was able to characterize my prototype circuit. The transfer function from a swept sine measurement looks very similar to the theoretically calculated transfer function (both of which are attached). The frequency response of the circuit was considered over the range 10 Hz - 10 kHz, which contains the eventual working range of the ISS (at least to my knowledge).

Note that OP27 op-amps were used instead of the high-speed AD829 op-amps that will be implemented in the actual design. This was done as a result of the limitations and inherent noise characteristics of the breadboard on which the prototype was built.

Unfortunately, I saved the wrong dataset (i.e. phase of the transfer function, not magnitude) and thus the presented function here is image generated by the SR785.

RXA: One must learn to use the python-GPIB interface to not lose data in the future.

Attachment 1: Prototype_Transfer_Function.png
Prototype_Transfer_Function.png
Attachment 2: Theoretical_Transfer_Function.png
Theoretical_Transfer_Function.png
  8109   Tue Feb 19 15:10:02 2013 JamieUpdateCDSc1iscex alive again

c1iscex is back up.  It is communicating with it's IO chassis, and all of it's models (c1x01, c1scx, c1spx) are running again.

The problem was that the IO chassis had no connection to the computer.  The One Stop card in the IO chassis, which is the PCIe bridge from the front-end machine and the IO chassis, was showing four red lights instead of the dozen or so green lights that it usually shows.  Upon closer inspection, the card appeared to be complaining that it had no connection to the host card in the front-end machine.  Un-illuminated lights on the host card seemed to be pointing to the same thing.

There are two connector slots on the expansion card, presumably for a daisy chain situation.  Looking at other IO chassis in the lab I determined that the cable from the front-end machine was plugged into the wrong slot in the One Stop card.  wtf.

Did someone unplug the cable connecting c1iscex to it's IO chassis, and then replug it in in the wrong slot?  A human must have done this.

  8108   Tue Feb 19 12:02:00 2013 JamieUpdateIOOIMC table levelling.

In order to address the issue of low MC1 OSEM voltages, Yuta and I looked at the IMC table levelling.  Looking with the bubble level, Yuta confirmed that the table was indeed out of level in the direction that would cause MC1 to move closer to it's cage, and therefore lower it's OSEM voltages.  Looking at the trends, it looks like the table was not well levelled after TT1 installation.  We should have been more careful, and we should have looked at the MC1/3 voltages after levelling.

Yuta moved weights around on the table to recover level with the bubble level.  Unfortunately this did not bring us back to good MC1 voltages.  We speculate that the table was maybe not perfectly level to begin with.  We decided to try to recover the MC1 OSEM voltages, rather than go solely with the bubble level, since we believe that the MC suspensions should be a good reference.  Yuta then moved weights around until we got the MC1/3 voltages back into an acceptable range.  The voltages are still not perfect, but I believe that they're acceptable.

The result is that, according to the bubble level, the IMC table is low towards MC2.  We are measuring spot positions now.  If the spot positions look ok, then I think we can live with this amount of skew.  Otherwise, we'll have to physically adjust the MC1 OSEMS.

Screenshot-Untitled_Window.png

  8107   Tue Feb 19 09:37:23 2013 SteveUpdateIOOlow MC1 OSEM voltage

 

    See TT DB25 pin swapping   elog#7869 

Attachment 1: MC1_MC3.png
MC1_MC3.png
Attachment 2: MC1_MC3_.png
MC1_MC3_.png
  8106   Tue Feb 19 08:42:31 2013 KojiUpdateVACETMX door open

[Steve, Yuta, Koji]

The ETMX heavy door was removed.

  8105   Tue Feb 19 08:06:02 2013 yutaUpdateElectronicsPOP path set up but AS55 is broken

I didn't use LED flash light. We learned from the past (elog #7355). I checked that POP55 and REFL55/165/33/11 are clearly responding to flash flight, but I didn't expect that much difference in DC gain.
I wonder why we could align AS beam to AS55 in Feb 8 (elog #8030), but not in Feb 15 (elog #8091). I will check during the pump down.

Quote:

10010 Ohm for POP55 vs 50 Ohm for AS55 (cf. http://nodus.ligo.caltech.edu:8080/40m/4763)

I wonder if you used an LED flash light, which emits no IR.

  8104   Tue Feb 19 05:42:28 2013 KojiUpdateElectronicsPOP path set up but AS55 is broken

10010 Ohm for POP55 vs 50 Ohm for AS55 (cf. http://nodus.ligo.caltech.edu:8080/40m/4763)

I wonder if you used an LED flash light, which emits no IR.

  8103   Tue Feb 19 02:23:40 2013 yutaBureaucracyGeneralaction items for PRFPMI

These are things need to be done for demonstrating PRFPMI using ALS.
All of these should be done before March 8!

CDS:
    - Fix c1iscex -JAMIE (done Feb 19: elog #8109)
    - Is ASS and A2L working? -JENNE
    - Are all whitening filters for PDs toggling correctly? -JENNE, JAMIE

PRMI locking:
    - Adjust I/Q rotation angles for error signals -JENNE, YUTA
    - Adjust filters -JENNE, YUTA
    - Coil balancing for BS (and ITMs/ETMs) -YUTA

PRC characterization in PRMI:

    - Measure PR2 loss from flipping -MANASA
    - Measure mode matching ratio -JENNE, YUTA
    - Measure finesse, PR gain -JENNE, YUTA
    - Calibrate PRM and/or ITM oplevs -MANASA, YUTA
    - Measure g-factor by tilting PRM or ITMs -JAMIE, YUTA
    - Calculate expected mode matching ratio and g-factor -JAMIE
    - Calculate expected finesse, PR gain -JENNE
    - Align aux laser into AS port? -ANNALISA?

ALS:
    - What's the end green situation? Optical layout changed? Laser temperature in CDS? -MANASA
    - What's the PSL green situation? Green trans cameras/PD? -JENNE, MANASA
    - Make ALS handing off to DARM/CARM LSC script -JENNE, YUTA
    - Demonstrate FPMI using ALS -JENNE, YUTA
    - Phase tracker characterization -YUTA, KOJI

PRFPMI:
    - Measure mode matching between PRC and arms -JENNE, YUTA
    - Measure PR gain -JENNE, YUTA
    - Calculate expected finesse, PR gain -JENNE

Others:
    - Update optical layout CAD after PR2 flipping -JAMIE, MANASA
    - AS55 situation? -YUTA
    - Look into PMC drift -JENNE, MANASA
    - Measure RFAM contribution to error signals -YUTA

 


Please fix, add or update if you notice anything.

  8102   Tue Feb 19 00:21:09 2013 yutaUpdateElectronicsPOP path set up but AS55 is broken

Hmm......
I thought AS55 is broken because it was not responding to the AS beam nor flashlight in DC. What's the DC gain difference between AS55 and POP55 (or REFL55)?

Quote:

I undertook the investigation of the AS55 PD. I found the PD is not broken.

  8101   Mon Feb 18 23:41:15 2013 SendhilUpdateAlignmentAligned IPPOS, IPANG and OPLEVS

[Yuta, Sendhil]

We aligned IPPOS, IPANG and all OPLEVs (except for ETMX and SRM).

 

1. First aligned the IPPOS by tweeking the steering mirrors inside the BS chamber.

2. Aligned the IPANG by tweeking the steering mirrors inside the BS chamber and ETMY chamber.

3. Aligned the OPLEVS for the BS and PRM was done by tweeking the steering mirrors inside the BS chamber and checked that OPLEV beams were not clipped. 

4. Centred the OPLEV beams for the ITMY and ETMY.

5. For the OPLEV of ITMX the alignment was done by tweeking the steering mirrors inside the ITMX chamber.

 

Attachment 1: IPANGIPPOSoplevs.png
IPANGIPPOSoplevs.png
  8100   Mon Feb 18 21:43:05 2013 KojiUpdateAlignmentPOP path set up but AS55 is broken

I undertook the investigation of the AS55 PD. I found the PD is not broken.

I tested the PD on the PD test bench and it works just fine.

I attatched the characterization result as there has been no detailed investigation of this PD as far as I remember.

The transimpedance gain at 55MHz is 420Ohm, and the shotnoise intercept current is 4.3mA.

Attachment 1: AS55.pdf
AS55.pdf AS55.pdf AS55.pdf
  8099   Mon Feb 18 18:28:01 2013 yutaBureaucracyAlignmentwe are going to pump down

We will start preparing for pumping down. Main goal for this is to demonstrate PRFPMI using ALS.
Here are to-dos before we pump down.

Feb 18 eveing
- check input beam and Y arm alignment again
- IPPOS/IPANG alignment
- check all oplevs

Feb 19 morning
- open ETMX chamber heavy door
- align BS to X end
- adjust OSEM values (added by YM)
- center beam on all AS optics
- make sure AS/REFL is clear
- take picture of flipped PR2 (added by YM)
- make sure green is not clipped by new PRM oplev mirrors  (added by YM)
- center all oplevs

Feb 19 afternoon - Feb 20 morning
- close PSL shutter
- close all heavy doors and put the access connector back
- start pumping down

Feb 20 evening
- start aligning IFO

  8098   Mon Feb 18 11:54:15 2013 Max HortonUpdateSummary PagesTiming Issues and Calendars

Crontab: The bug of data only plotting until 5PM is being investigated.  The crontab's final call to the summary page generator was at 5PM.  This means that the data plots were not being generated after 5PM, so clearly they never contained data from after 5PM.  In fact, the original crontab reads:

0 11,5,11,17 * * * /users/public_html/40m-summary/bin/c1_summary_page.sh 2>&1

I'm not exactly sure what inspired these entries.  The 11,5,11,17 entries are supposed to be the hours at which the program is run.  Why is it run twice at 11?  I assume it was just a typo or something.

The final call time was changed to 11:59PM in an attempt to plot the entire day's data, but this method didn't appear to work because the program would still be running past midnight, which was apparently inhibiting its functionality (most likely, the day change was affecting how the data is fetched).  The best solution is probably to just wait until the next day, then call the summary page generator on the previous day's data.  This will be implemented soon.

Calendars: Although the calendar tabs on the left side of the page were fixed, the calendars displayed at: https://nodus.ligo.caltech.edu:30889/40m-summary/calendar/ appear to still have squished together text.  The calendar is being fetched from https://nodus.ligo.caltech.edu:30889/40m-summary/calendar/calendar.html and displayed in the page.  This error is peculiar because the URL from which the calendar is being fetched does NOT have squished together text, but the resulting calendar at 40m-summary/calendar/ will not display spaces between the text.  This issue is still being investigated.

  8097   Mon Feb 18 00:03:46 2013 ZachUpdateComputer Scripts / ProgramsARBCAV v3.0

I have uploaded ARBCAV v3.0 to the SVN. The major change in this release, as I mentioned, is the input/output handling. The input and output are now contained in a single 'model' structure. To define the cavity, you fill in the substructure 'model.in' (e.g., model.in.T = [0.01 10e-6 0.01]; etc.) and call the function as:

model = arbcav(model);

Note: the old syntax is maintained as legacy for back-compatibility, and the function automatically creates a ".in" substructure in the output, so that the user can still use the single-line calling, which can be convenient. Then, any individual parameter can be changed by changing the appropriate field, and the function can be rerun using the new, simpler syntax from then on.

The function then somewhat intelligently decides what to compute based on what information you give it. Using a simple option string as a second argument, you can choose what you want plotted (or not) when you call. Alternatively, you can program the desired functionality into a sub-substructure 'model.in.funct'.

The outputs are created as substructures of the output object. Here is an example:

 

>> th = 0.5*acos(266/271) *180 /pi;

OMC.in.theta = [-th -th th th];

OMC.in.L = [0.266 0.284 0.275 0.271];

OMC.in.RoC = [1e10 2 1e10 2];

OMC.in.lambda = 1064e-9;

OMC.in.T = 1e-6 * [8368 25 8297 33];

OMC.in.f_mod = 24.5e6;

>> OMC

OMC = 

    in: [1x1 struct]

>> OMC = arbcav(OMC,'noplot')

Warning: No loss given--assuming lossless mirrors 

> In arbcav at 274 

OMC = 

         in: [1x1 struct]

        FSR: 2.7353e+08

        Lrt: 1.0960

    finesse: 374.1568

    buildup: 119.6956

         df: [1000x1 double]

      coefs: [1000x4 double]

        HOM: [1x1 struct]

>> OMC.HOM

ans = 

      f: [1x1 struct]

    pwr: [1x1 struct]

>> OMC.HOM.pwr

ans = 

    carr: [15x15 double]

     SBp: [15x15 double]

     SBm: [15x15 double]

 

Some other notes:

  • The annoying Mdo.m has been internalized; it is no longer needed.
  • For the next release, I am working on including:
    • Finite mirror thickness/intracavity refractive elements - If, for god knows what reason, you decide to put a mirror substrate within a cavity 
    • Mode overlap - Calculating the overlap of an input beam to the cavity
    • Mode matching - Calculating a mode matching telescope into the cavity for some defined input beam
    • Anything else?

I have added lots of information to the help header, so check there for more details. As always, your feedback is greatly appreciated.

  8096   Sun Feb 17 19:27:19 2013 ranaUpdateSUSPRM coil balanced

 I will check out the AS55 situation tomorrow. Just put it on my desk.

MC Autolocker was disabled - I enabled it.

For the F2P.py, you should look at how we did this with the script written 8 years ago in csh. There we stored the initial values in a file (so they don't get blow away if someone does CTRL-C). Your python script should have a trap for SIGINT so that it dies gracefully by restoring the initial values. In order to have the smooth value adjustment, you must first set the TRAMP field for all the coil gains to 2 and then switch. Make sure that the lockin ignores the first few seconds of data after making this switch or else it will be hugely biased by this transient.

For the PRM OL use as a F2A reference, you also have to take into account that the OL beam is hitting the PRM surface at non-normal incidence. IF it is a large angle, there will be a systematic error in the setting of the F2Y values.

  8095   Sat Feb 16 19:23:17 2013 yutaUpdateLSCPR2 flipped PRMI locked

It is my pleasure to announce that the first lock of PR2 flipped PRMI was succeeded.



POP looks very nice. TEM00 and not wobbling.
We need more I/Q phase and gain/filter adjustment and characterization soon.

Some more details:
  MICH error signal: AS55_Q_ERR (using POP55 PD; phase rotation angle 70 deg)
  PRCL error signal: REFL11_I_ERR (phase rotation angle 80 deg)
  MICH feedback: BS (MICH_GAIN = -60)
  PRCL feedback: PRM (PRCL_GAIN = -0.5)

  8094   Sat Feb 16 18:32:01 2013 yutaSummaryRF Systemphase tracker: OLTF

I measured openloop transfer function of the phase tracking loop for the first characterization of phase tracker.

What is phase tracker:

  See elog #6832.
  For ALS, we use delay-line frequency discriminator, but it has trade-off between sensitivity and linear range. We solved this trade-off by tacking the phase of I/Q signals.
  Figure below is the current diagram of the frequency discriminator using phase tracker.
phasetracker.png

OLTF of phase tracking loop:
  Below. UGF at 1.2 kHz, phase margin 63 deg for both BEATX and BEATY. Phase delay can be clearly explained by 61 usec delay. This delay is 1 step in 16 KHz system.
  Note that UGF depends on the amplitude of the RF input. I think this should be fixed by calculating the amplitude from I/Q signals.
  BEAT(X|Y)_PHASE_GAIN were set to 300, and I put -3dBm 100 MHz RF signal to the beatbox during the measurement.
BEATX:BEATXFINEPHASE_OLTF.png BEATY:BEATYFINEPHASE_OLTF.png

Other measurements needed:

 - Linear range: By sweeping the RF input frequency and see sensitivity dependence.
 - Bandwidth: By measuring transfer function from the modulation frequency of the RF input to phase tracker output.
 - Maximum sensitivity: Sensitivity dependence on delay-line length (see PSL_Lab #825).
 - Noise: Lock oscillator frequency with phase tracker and measure out-of-loop frequency noise with phase tracker.
 - Sensitivity to amplitude fluctuation: Modulate RF input amplitude and measure the sensitivity.

  8093   Sat Feb 16 17:27:26 2013 yutaUpdateSUSPRM coil balanced

PRM coil gains and f2a filters are adjusted for PRMI work.
It seems like UR/LL coil gains were ~10 % larger than others, and f2a filters changed by few %.

What I did:
  1. Tried to lock PRMI but when I turn on PRCL lock, PRM reflection looked like it tends to go up and left in REFL camera (last night).

  2. So, I set up PRM oplev back, by steering PRM oplev mirrors on the BS table (last night).

  3. Turned PRM oplev sero on, f2a filters off, and ran

> /opt/rtcds/caltech/c1/scripts/SUS/F2P_LOCKIN.py -o PRM

  I had to fix F2P_LOCKIN.py because it assumed some OUTPUT buttons in LOCKIN1 filters to be ON.
  Also, I had to restore filters in LOCKIN1 (8.5 Hz bandpass filter etc.) because their names were changed. To do this, I copied filters needed from /opt/rtcds/caltech/c1/chans/filter_archive/c1sus/C1SUS_110916_162512.txt, renamed LOCKIN1_(I|Q|SIG) with LOCKIN1_DEMOD_(I|Q|SIG), and pasted to the current filter bank file. I checked that they look OK with foton after editing the file.

  This measurement takes about 30 minutes. I ran several times to check consistency. There was ~ 0.1 % standard deviation for the measurement results.

  4. By putting measured coupling coefficients and PRM pendulum frequency (f0=0.993 Hz) to /opt/rtcds/caltech/c1/scripts/SUS/F2Pcalc.py, I got new f2a filters.

  5. Overwrote f2a filters in C1:SUS-PRM_TO_COIL_(1-4)_1 FM1 with new ones, and turned  new f2a filters on.

Result:
  Below is the DC gain adjustment result from F2P_LOCKIN.py;

multiplier factors are :
UL = 1.141525
UR = 0.879997
LR = 1.117484
LL = 0.860995
Set C1:SUS-PRM_ULCOIL_GAIN to 1.04990177238
Set C1:SUS-PRM_URCOIL_GAIN to -0.983396190716
Set C1:SUS-PRM_LRCOIL_GAIN to 0.954304254663
Set C1:SUS-PRM_LLCOIL_GAIN to -0.971356852259


  So, UR/LL coil gains somehow got ~10 % larger than other two since last coil balancing.

  Measured coupling coefficients from F2P_LOCKIN.py were

- measured coupling coefficients are :
P2P(POS=>PIT) = 0.014993
P2Y(POS=>YAW) = 0.001363


  New f2a filters are plotted below. They look fairly different compared with previous ones.
PRM_f2a.png


 

We need better F2P_LOCKIN.py:
  Some one should make F2P_LOCKIN.py better. The main problem is the sudden gain change when starting diagonalization at low frequency. It sometimes trips off the watchdog.

Some elogs related:
  Kiwamu made f2a filters in Sep 2011: elog #5417
  Koji adjusting DC gains in Jan 2013: elog #7969

  8092   Fri Feb 15 21:22:29 2013 yutaUpdateElectronicsAS55 replaced with POP55 PD

I temporarily replaced AS55 PD with PD labeled "POP55(POY55)".
I think POP55 is working because I could lock MI with this PD using AS55_Q_ERR as an error signal. I rotated I/Q phase (C1:LSC-AS55_PHASE_R) to 70 deg by minimizing ASDC during MI lock.

POP55 PD was freely sitting on the ITMX table.
I will leave AS55 PD at free space of the AP table. Someone, please look into it.

  8091   Fri Feb 15 20:07:28 2013 yutaUpdateAlignmentPOP path set up but AS55 is broken

[Manasa, Yuta]

We set up POP camera and POPDC PD, and centered REFL PDs.
We also tried to center AS55 PD, but AS55 seems to be broken.

What we did:
 1. POP path alignment:
   Shot green laser pointer from ITMX table at where POPDC PD was sitting and centered green beam at optics in the POP path. Steered POPM1/M2 mirrors in the ITMX chamber to make green laser overlap with the PRM-PR2 beam as far as I can reach from ITMX chamber. We removed some ND filters and a BS for attenuating POP beam because POP power was somehow so low. Currently, POP is pick-off of the beam which goes from PRM to PR2.

 2. POP camera and PD:
   We first used camera to find the beam at where POPDC PD was sitting because it is much easier to find focused beam. Put an iris in front of the camera, and put POP DC behind it. Steered a mirror in front of PD to maximize DC output.

 3. REFL PDs:
   Steered mirrors in the REFL path to center the beam and maximized DC outputs, as usual.

 4. AS55:
   AS55 was not responding very much to the flashlight nor AS beam. C1:LSC-ASDC_OUT looked funny. By swapping the ribbon cables of AS55, REFL55, and REFL165, I confirmed that AS55 PD itself is broken. Not the ribbon cable nor PD circuit at LSC rack. I don't know what happened. AS55 was working on Feb 8 (elog #8030).

Result:
  We aligned PRMI coarsely. POP(right above) looks much better than before. REFL (left below) still looks elliptic, but ellipticity differs with the position on the camera. Some astigmatism is happening somewhere. AS (right below) looks pretty nice with MI aligned.


Next:
  1. Fix AS55? Or replace it with POP55 PD, which is currently unused.
  2. Confirm we are getting the right error signals or not, and lock PRMI.

  8090   Fri Feb 15 17:11:13 2013 ChloeUpdate QPD circuit pictures

 I took better pictures of the circuits of the QPD and spent a couple of hours with a multimeter trying to figure out how all the connections worked. I will continue to do so and analyze the circuits over the weekend to try to understand what is going on. I also have an old SURF report that Eric sent me that is similar to the design I was planning to use to sum the pitch and yaw signals. I will try and look at this over the weekend. 

Attachment 1: IMG_0337.JPG
IMG_0337.JPG
Attachment 2: IMG_0338.JPG
IMG_0338.JPG
  8089   Fri Feb 15 16:09:19 2013 KojiUpdateGeneralbeam traps ready to be installed

For the hexagonal one, insert one of the glass plate only half. Use a 1"x.5" piece if exists.

For the diamond one, you don't need the forth glass piece.

 

Attachment 1: HexBeamDump.pdf
HexBeamDump.pdf
Attachment 2: DiamondBeamDump2in.pdf
DiamondBeamDump2in.pdf
  8088   Fri Feb 15 15:21:07 2013 JamieUpdateComputersc1iscex IO-chassis dead

I appears that the c1iscex IO-chassis is either dead or in a very bad state.  The PCIe interface card in the IO-chassis is showing four red lights, where it's supposed to be showing a dozen or so green lights.  Obviously this is going to prevent anything from running.

We've had power issues with this chassis before, so possibly that's what we're running into now.  I'll pull the chassis and diagnose asap.

 

  8087   Fri Feb 15 09:58:49 2013 SteveUpdateGeneralbeam traps ready to be installed

Black-green glass traps are ready for light in vacuum. I can assemble more if needed. These three sizes are available.

Attachment 1: IMG_0083.JPG
IMG_0083.JPG
Attachment 2: IMG_0084.JPG
IMG_0084.JPG
  8086   Fri Feb 15 01:51:43 2013 JenneSummaryAlignmentIFO aligned and ready for PRMI locking

 

Yuta and Manasa, you guys are awesome!

Small, inconsequential point:  The camera image in the upper right of your video is the *back* of the Faraday in our usual nomenclature.  The camera is listed in the videoswitch script as "FI_BACK".  The camera looking at the "front" of the Faraday is just called "FI". 

  8085   Fri Feb 15 01:41:02 2013 Manasa, YutaSummaryAlignmentIFO aligned and ready for PRMI locking

[Yuta, Manasa, Jenne, Jamie, Steve]

IFO aligned and ready for PRMI locking

Alignment procedure

0. Measured MC centering (off by 5mrad) before getting the doors off.

1. Got the TTs to 0.0 in pitch and yaw.

2. Using the MMTs, the beam was centered on the TTs.

3. TT1 was adjusted such that the incident beam was centered at PRM (with target).

4. TT2 was adjusted such that the beam passed through the center of BS (with target).

5. Centered the beam on PR2 by sliding it on the table.

6. Moved PR2 and tweaked TT2 to center the beam on ITMY and BS respectively.

7. Using TTs, we got the beam centered on ETMY while still checking the centering on ITMY.

8. ITMY was adjusted such that it retro-reflected at the BS.

9. ETMY was aligned to get a few bounces in the arm cavity.

10. Centered on ITMX by adjusting BS and then tweaked ITMX such that we retro-reflected at BS.

11. At this point we were able to see the MI fringes at the AS port.

12. Tweaked ITMX to obtain reflected MI fringes in front of MMT2.

13. By fine adjustments of the ITMs, we were able to get the reflected MI to go through the faraday  while still checking that we were retro-reflecting at the BS.

14. Tweaked the PRM, such that the PRM reflected beam which was already visible on the 'front face back face of faraday' camera went through the faraday and made fine adjustments to see it fringing with the reflected MI that was already aligned.

15. At this point we saw the REFL (flashing PRMI) coming out of vacuum unclipped and on the camera.

16. Started with alignment to get the AS beam out of vacuum. We tweaked OM1 and OM2 (steering mirrors in the ITMY chamber) to center the beams on OM4 and OM3 (steering mirrors in the BSC) respectively.

17. We then adjusted steering mirrors OM5 and OM6 (in the OMC chamber) such that the beam went unclipped out of vacuum.

18. Note that we took out the last steering mirror (on the AS table) in front of the AS camera, so that we can find the AS beam easily. This can be fixed after we pump down.

 

 

Tomorrow

  0. REFL still looks like an egg, but leave it .

  1. Align PRMI (no more in-vac!) .

  2. Align POP/REFL/AS cameras and PDs.

  3. Setup PRM/BS/ITMX/ITMY oplevs.

  4. Balance the coils on these mirrors.

  5. Lock PRMI.

  8084   Thu Feb 14 10:42:41 2013 JamieSummaryAlignmentMMT, curved TTs does not explain beam ellipticity at Faraday

After using alamode to calculate the round-trip mode of the beam at the Faraday exit after retro-reflection form the PRM, I'm not able to blame the MMT and TT curvature for the beam ellipticity.

I assume an input waist at the mode cleaner of [0.00159, 0.00151] (in [T, S]).  Propagating this through the MMT to PRM, then retro-reflecting back with flat TTs I get

w_t/w_s = 0.9955,  e = 0.0045

If I give the TTs a -600 m curvature, I get:

w_t/w_s = 1.0419,  e = 0.0402

That's just a 4% ellipticity, which is certainly less than we see.  I would have to crank up the TT curvature to -100m or so to see an ellipticity of 20%.  We're seeing something that looks bigger than 50% to me.

Below are beam size through MMT + PRM retro-reflection, TT RoC = -600m:

 

T.pdfS.pdf

  8083   Thu Feb 14 08:29:41 2013 SteveUpdateIOOlow MC1 OSEM voltage

MC1 -  LR, LL, UL & UR  OSEMs should be adjusted to get  1.2V

Attachment 1: Feb14_2013.png
Feb14_2013.png
  8082   Thu Feb 14 00:10:12 2013 yutaSummaryAlignmentREFL is not clipped

Let's wait for astigmatism calculation.
In either case(clipping or astigmatism), it takes time to fix it. And we don't need to fix it because we can still get LSC signal from REFL.
So why don't we start aligning input TTs and PRMI tomorrow morning.

Take the same alignment procedure we did yesterday, but we should better check REFL more carefully during the alingment. Also, use X arm (ETMX camera) to align BS. We also have to fix AS steering mirrors in vacuum. I don't think it is a good idea to touch PR2 this time, because we don't want to destroy sensitive PR2 posture.


Calculations need to be done in in-air PRMI work:
  1. Explanation for REFL astigmatism by input TTs (Do we have TT RoCs?).
  2. Expected g-factor of PRC (DONE - elog #8068)
  3. What's the g-factor requirement(upper limit)?
    Can we make intra-cavity power fluctuation requirement and then use PRM/2/3 angular motion to break down it into g-factor requirement?
    But I think if we can lock PRMI for 2 hours, it's ok, maybe.
  4. How to measure the g-factor?
    To use tilt-and-measure-power-reduction method, we need to know RoC of the mirror you tilted. If we can prove that measured g-factor is smaller than the requirement, it's nice. We can calculate required error for the g-factor measurement.

  8081   Wed Feb 13 22:09:26 2013 JenneUpdateAlignmentREFL is not clipped

We need to calculate whether this level of astigmatism is expected from the new active TT mirrors, but I claim that the beam is not clipped.

As proof, I provide a video (PS, why did it take me so long to be converted to using video capture??).  I'm just showing the REFL camera, so the REFL beam as seen out on the AS table.  I am moving PRM only.  I can move lots in pitch before I start clipping anywhere.  I have less range in yaw, but I still have space to move around.  This is not how a clipped beam behaves.  The clipping that I see after moving a ways is coincident with clipping seen by the camera looking at the back of the Faraday.  i.e. the first clipping that happens is at the aperture of the Faraday, as the REFL beam enters the FI.  

Also, I'm no longer convinced entirely that the beam entering the Faraday is a nice circle.  I didn't check that very carefully earlier, so I'd like to re-look at the return beam coming from TT1, when the PRM is misaligned such that the return beam is not overlapped with the input beam.  If the beam was circular going into the Faraday, I should have as much range in yaw as I do in pitch.  You can see in the movie that this isn't true.  I'm voting with the "astigmatism caused by non-flat active TT mirrors" camp. 

  8080   Wed Feb 13 19:41:07 2013 yutaSummaryGeneralpossible explanations to oval REFL beam

We checked that REFL beam is already oval in the vacuum. We also centered in-air optics, including lens, in the REFL path, but REFL still looks bad.

By using IR card in vacuum, PRM reflected beam looks OK at MMTs and at the back face of the Faraday. But the beam looks bad after the output aperture of the Faraday.

  8079   Wed Feb 13 19:30:45 2013 KojiSummaryGeneralpossible explanations to oval REFL beam

>> "What has changed since:"

Recently the REFL path has been rearranged after I touched it just before Thanksgiving.
(This entry)

If the lenses on the optical table is way too much tilted, this astigmatism happens.
This is frequently observed as you can find it on the POP path right now.

Also the beam could be off-centered on the lens.

I am not sure the astigmatism is added on the in-air table, but just in case
you should check the table before you put much effort to the in-vacuum work.

  8078   Wed Feb 13 19:09:32 2013 yutaSummaryGeneralpossible explanations to oval REFL beam

[Jenne, Manasa, Jamie, Yuta]

The shape of the REFL beam reflected from PRM is oval after the Faraday.
We tried to fix it by MC spot position centering and by tweaking input TT1/TT2/PRM. But REFL still looks bad (below).

REFL_1044844506.bmp

What has changed since:
  REFL looks OK in mid-Dec 2012. Possibly related things changed are;

  1. New active input TTs with new mirrors installed
  2. Leveling of IMC stack changed a little (although leveling was done after installing TTs)

Possible explanations to oval REFL:
  A. Angled input beam:
    Input beam is angled compared with the Faraday apertures. So, beam coming back from PRM is angled, and clipped by the Faraday aperture at the rejection port.

  B. Mode mis-match to PRM:
    New input TTs have different curvatures compared with before. Input mode matching to PRM is not good and beam reflected from PRM is expanding. So, there's clipping at the Faraday.

  C. Not clipping, but astigmatism:
    New input TTs are not flat. Incident angle to TT2 is ~ 45 deg. So, it is natural to have different tangential/sagittal waist sizes at REFL.

How to check:
  A. Angled input beam:
    Look beam position at the Faraday apertures. If it doesn't look centered, the incident beam may be angled.
   (But MC centering didn't help much......)

  B. Mode mis-match to PRM:
    Calculate how much the beam size will be at the Faraday when the beam is reflected back from PRM. Put some real numbers to curvatures of input TTs for calculation.

  C. Not clipping, but astigmatism:
    Same calculation as B. Let's see if REFL is with in our expectation or not by calculating the ratio of tangential/sagittal waist sizes at REFL.

  8077   Wed Feb 13 16:31:08 2013 JenneUpdatePSLPMC pitch input tuned, MC yaw input tuned

[Jenne, Yuta]

I looked at PMCR camera on the MC1 tv, and tweaked up the beam going into the PMC - it only needed a little bit of pitch.

Yuta and I measured the MC spots, determined (consistent with my measurements this morning) that they were only off in yaw.  We touched the 2nd steering mirror in the zigzag on the PSL table in yaw a small amount (top of knob away from me), realigned the MC, and things were good.  The plot is zoomed in to show only measurements taken today.  2 in the morning, before anything in the IFO room was touched.  1 this afternoon after tweaking PMC.  1st attempt at PSL beam tweaking was successful, 2nd measurement confirms it wasn't a fluke.

MCspots_13Feb2013.png

  8076   Wed Feb 13 14:21:19 2013 JenneUpdateLockingPRC cavity gains

Quote:

With 1500ppm loss on both PR2 and PR3, 150ppm arm cavity loss:

We get a PRC gain for the CARRIER (non-flipped folding) of 21, and PRC gain (flipped folding) of 20.  This is a 4.7% loss of carrier buildup.

We get a PRC gain for the SIDEBANDS (non-flipped folding) of 69, and PRC gain (flipped folding) of 62.  This is an 8.8% loss of sideband buildup.

 With a PR2 loss of 896ppm (from the plot on the wiki), but no loss from PR3 because we didn't flip it, and the same 150ppm round trip arm cavity loss, I get:

Carrier gain = 21.0

Sideband gain = 66.7

(No loss case, with an extra sig-fig, so you can see that the numbers are different:  Carrier = 21.4, Sideband = 68.8 .)

So, this is 1.6% loss of carrier buildup and 3.1% loss of sideband buildup.  Moral of the story - G&H's measured AR reflectivity is less than Rana's guess, and we didn't flip PR3, so we should have even less of a power recycling gain effect than previously calculated.

  8075   Wed Feb 13 09:28:56 2013 SteveUpdateOpticsG&H - HR plots

 

 Gooch & Housego optics order specification from 03-13-2010

Side 1: HR Reflectivity >99.99 % at 1064 nm for 0-45 degrees for S & P polarization

Side 2: AR coat R <0.15

The HR coating scans uploaded to 40mwiki / Aux optics today

  8074   Wed Feb 13 01:26:08 2013 yutaSummaryGeneralrough analysis of aligned PRM-PR2 mode scan

Koji was correct.

When you estimate the variance of the population, you have to use unbiased variance (not sample variance). So, the estimate to dx in the equations Koji wrote is

dx = sqrt(sum(xi-xavg)/(n-1))
   = stdev*sqrt(n/(n-1))


It is interesting because when n=2, statistical error of the averaged value will be the same as the standard deviation.

dXavg = dx/sqrt(n) = stdev/sqrt(n-1)

In most cases, I think you don't need 10 % precision for statistical error estimation (you should better do correlation analysis if you want to go further). You can simply use dx = stdev if n is sufficiently large (n > 6 from plot below).
unbiased.png



Quote:

Makes sense. I mixed up n and n-1

Probability function: X = (x1 + x2 + ... + xn)/n, where xi = xavg +/- dx

Xavg = xavg*n/n = xavg

dXavg^2 = n*dx^2/n^2
=> dXavg = dx/sqrt(n)

Xavg +/- dXavg = xavg +/- dx/sqrt(n)

 

  8073   Tue Feb 12 23:24:17 2013 yutaUpdateAlignmentIFO alignment in prep for in-air PRMI

[Manasa, Yuta]

Lot's of alignment work, still no AS beam. REFL is clipped by Faraday output aperture......
Our guess is that this is because
we skipped MC centering.

Alignment procedure we took:
 1. AM work: Aligned input beam using TT1/TT2
   such that the beam hits ETMY and ITMY at the center.

 2. Coarsely aligned ITMY
   such that the ITMY retro-reflected beam hits BS at the center.

 3. Aligned ETMY (we didn't actually move ITMY)
   such that Y arm flashes.
   This tells you that ITMY is aligned well to the incident beam.

 4. Aligned BS
   such that the beam hits ITMX at the center.

 5. Aligned ITMX
   such that the ITMX retro-reflected beam hits BS at the center.
   At this point, we saw MI fringes at AS port.

 6. Fine alignment of ITMX:
   MI reflected beam was not overlapping in front of BS after it was reflected by PRM.
   We used this longer REFL path to tune alignment of ITMX to ITMY reflected beam.
   We saw MI fringe at REFL port coming out of the chamber, but it was clipped.

 7. Aligned PRM
   by looking at REFL beam from PRM on the back face of Faraday (video FI_BACK).
   We fine tuned the alignment such that PRM retro-relfected beam hits BS at the center and REFL beam from PRM overlaps with the MI fringes at the back face of Faraday.

 8. Clipping of REFL at the Faraday output aperture:
   We confirmed that the shape of the REFL beam from PRM was OK at the back face of Faraday. But some how, it was clipped at the output aperture. So, REFL beam coming out of the chamber is clipped now.

 9. Tried to get AS beam out of the chamber:
   We tweaked steering mirrors after SRM to get AS beam out of the chamber. But, we lost the AS beam between the very last folding mirrors (OMPO and OM6) in the OMC chamber......


Discussion:
 1. Why clipping at the Faraday output aperture?
   In principle, if PRM reflects the incident beam at normal incidence, it should pass the Faraday unclipped. But it's not!
   Our guess is that the incident beam does not go well centered through the apertures of the Faraday. I think we have to do MC centering to get good pointing to the Faraday.
   We also see that MI fringe at the back face of the Faraday is at the edge of its aperture, after all of these alignment work (we even used Y arm!). This tells you that some thing is wrong.

 2. Why did you guys lose the AS beam?
   AS beam is too weak after reflecting off of OMPO. The beam was neither visible on IR cards nor IR viewers. The beam is weaker than usual because PMC transmission is ~0.7 and MC REFL is getting high (~ 0.7). We didn't want to realign MC after all of this work today.


Tomorrow (my suggestion):
  1. Align PMC (for higher power).
  2. MC centering.
  3. Input beam steering using TTs and redo the same alignment procedure (it shouldn't take longer than today).
      ==> Center beam on PR2  (Added by Manasa)
  4. Maybe we should better check PRM reflection at REFL port after the Faraday, before doing the full alignment work.
  5. Align AS, REFL, POP PDs/cameras.
  6. Setup PRM/BS/ITMX/ITMY oplevs.
  7. Balance the coils on these mirrors.
  8. Lock PRMI.


What needs to be done before pumping down:
  1. PRMI characterization: PR gain and g-factor
   How can we do the g-factor measurement? Use additional laser? Kakeru method (elog #1434; we need to calibrate mirror tilt to do this)?
  2. Glitch study in PRMI locking. If still glitchy, we have to do something. How is beam spot motion? (elog #6953)
  3. Fine alignment of the flipped PR2.
  4. Fine alignment of IFO using both arms.

ELOG V3.1.3-